Leche Fresca is a wholesaler that commercializes dairy products in the metropolitan area of Mexico City. The company serves both modern (convenience stores and supermarkets) and traditional (mom and pop stores) channels. There are two main differences in how Lecha Fresca attends these two markets: order processing and delivery.

In the case of traditional channel, sales representatives visit in-person the small mom and pop store to take the customer order. This is different for modern channel. In this case, a sales representative makes a call to the convenience store or supermarket to take a new order. Usually, 80% of sales representatives' weekly time is spent on serving traditional stores, while 20% is dedicated to modern stores. Additionally, 60% of IT equipment hours are used to process the orders from traditional channel and 40% from modern channel every week.

The delivery process is also different. Due to small drop sizes, the company trucks visit more frequently (2 or 3 times a week) the same mom and pop store. In contrast, a weekly shipment is fairly common among modern retailers. On average, 80% of drivers' weekly time is spent delivering products to traditional retailers (20% for modern stores). Similarly, 70% of fuel is consumed distributing the goods to traditional stores (30% to modern stores) each week.

On a weekly basis, 250 orders are received from 50 traditional retailers and 350 orders come from 18 modern stores.

Leche Fresca has identified the following relevant activities:

  • Order processing, traditional channel
  • Order processing, modern channel
  • Delivery, traditional channel
  • Delivery, modern channel

Part 1

As a first step to understand the cost to serve customers, you need to find out what the costs for the activities are. Leche fresca has identified the following relevant weekly costs:

  • Personnel (sales representative), order processing: 710
  • Other than personnel, order processing: 270
  • Personnel (drivers and helpers), delivery: 1280
  • Fuel and operating costs, delivery: 900

Based on the information above, what are the total cost that should be allocated to the following two activities?

Order processing for traditional channel: ?

Delivery for modern channel: ?

Part 2

Having calculated the total cost, you want to find the cost driver rate for each activity. After talking to the management of the company, you decide on the following cost drivers for the activities:

  • Order processing for traditional channel: Number orders
  • Order processing for modern channel:: Number of orders
  • Delivery for traditional channel: Number of customers
  • Delivery for modern channel: Number of customers

Based on the information above, what are the cost driver rates for the following two activities?

Order processing for modern channel: ?

Delivery for traditional channel: ?

Part 3

Having calculated the cost driver rates, you are interested in finding out what the cost-to-serve would be to attend new city areas. In particular, you are interested in the following two city areas:

Zone A Zone B
Number of orders (traditional channel) 20 12
Number of customers (traditional channel) 10 18
Number of orders (modern channel) 20 12
Number of customers (modern channel) 10 18

Based on your calculations, what is the cost to serve the two city zones?

Zone A

Zone B
Answer #1

There are basically 2 types of stores and 2 types of task. Thus the combination of them provides 4 options.

Part 1

Traditional stores/channels:

We need to handle 250 orders from 50 traditional stores. (Assuming that 250 is the number of total orders from 50 customers)

This means the order processing cost associated with traditional channel is

250*(710+270) = 245000 per week

Modern channel

We need to handle 350 orders from 18 stores.

This means the delivery cost for modern channel is

350*(1280+900) = 763000 per week

Part 2

Order processing cost for modern channel is 350*(710+270) = 343000

Delivery cost for traditional channel is 250*(1280+900) = 545000

We have 50 traditional stores and 18 modern stores with orders of 250 and 350 respectively.

Cost driver rates

Order processing for modern channel = order processing cost/number of orders = 343000/350 = 980

Delivery for traditional channel = delivery cost/number of customers = 545000/50 = 10900

Part 3

The calculated values based on the rates and the provided information is shown below

Zone A Zone B
# of orders (TC) 20 12
# of customers (TC) 10 18
# of orders (MC) 20 12
# of customer (MC) 10 18
Rates Cost - Zone A Cost - Zone B
Order processing (TC) 980 19600 11760
Order processing (MC) 980 19600 11760
Delivery (TC) 10900 109000 196200
Delivery (MC) 42388.89 423888.8889 763000
Total cost 572088.8889 982720